Consider an individual who earns $100,000, pays $7,000 in child care expenses for one child, pays the maximum ($26,010) in Registered Pension Plan (RPP) contributions, makes charitable donations of $200 (15 percent credit), and pays $20,000 in medical expenses. Medical expenses in excess of 3 percent of net income are deductible at 20 percent. The statutory marginal tax rate is 17 percent.
What is the federal tax payable? What is the individual’s actual (effective) marginal tax rate?
1
Expert's answer
2019-06-12T11:11:09-0400

the federal tax payable = ( 100000 - 7000 - 26010 - ( 200 - 200*15/100) - 20000 *20/100 ) *17/100 = 62820 *17/100 = 10679.4

The effective tax rate is the average tax rate paid by a corporation or an individual. 

An individual's effective tax rate is calculated by dividing the "Total Tax" on "Taxable Income."

An individual's effective tax rate = 10679.4 / 62820 = 0.17